Solve the following pairs simultaneous equations.
4x - y = 12
2y - 8 = -24

Answer by Alan3354(69443)

Solve the following pairs simultaneous equations.
4x - y = 12
2y - 8 = -24 --> y = -8
-------
4x - y = 12
4x + 8 = 12
x = 1